OpenStudy (anonymous):

if after factoring a quadratic equation I end up with something like this: N(N-2)=72 How can I find the roots? Please help

OpenStudy (anonymous):

n^2-2n+1=73 (n-1)^2=73 n-1= + or - (73)^(1/2) n=+ or - (73)^(1/2) +1

OpenStudy (anonymous):

This method is called completing the square, and works with all quadratic forms.

OpenStudy (anonymous):

\[n^2 - 2n = 73\] \[use\ the\ sridharan's\ formula,\ i.e\ x=1\pm 2\ sqrt{74}\]

OpenStudy (anonymous):

Sorry to note, but the above answer did not properly transcribe the original equation, thus yielding an incorrect solution.

OpenStudy (anonymous):

x1=1+sqrt73 ans x2=1-sqrt73 answ

OpenStudy (anonymous):

did you get it emunrra?

OpenStudy (anonymous):

No really

OpenStudy (anonymous):

use the quadratic formula in these eq ax^2 + bx +c x=(-b-+sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

N(N-2)=72 is already factored. You dont have to expand it (that is what i think) so for me x1 and x2 must be: x1=72 x2=74 Not sure of that

OpenStudy (anonymous):

N(N-2)=72 N^2 -2N =72 N^2 -2N -72=0 USE a=1, b=-2,c=-72

OpenStudy (anonymous):

N^2 -2N -72=0 this is also x^2 -2x -72=0 USE a=1, b=-2,c=-72

OpenStudy (anonymous):

You do not need to resort to the formula. You can solve this problem without it, using only fundamental arithmetic operations. \[N(N-2)=72\] \[N^2-2N=72\] \[N^2-2N+1=73\] \[(N-1)^2=73\] \[N-1=\pm\sqrt{73}\] \[N= \pm\sqrt{73}\ +1\]

OpenStudy (anonymous):

ahhh ok. Got it.

OpenStudy (anonymous):

correct, if you use factoring

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

yw emunrra...lol..good luck...